OpenStudy (anonymous):

how do i solve square root of -54 + the square root of -96 ?

OpenStudy (mathmate):

\[\sqrt{-54}+\sqrt{-96}\]\[=i \sqrt{9.6} + i \sqrt{16.6}\]\[=3i \sqrt6 + 4i \sqrt6\]\[=7i \sqrt6\]

OpenStudy (anonymous):

why cant it be \[-7\sqrt{6}\]?

OpenStudy (mathmate):

It depends on the question. \[-\sqrt{54}-\sqrt{96}\] would give \( -7 \sqrt{6} \).

OpenStudy (mathmate):

The square-root of a negative quantity involves i where \( i^2 = -1\).

OpenStudy (anonymous):

ok then your right. i was just making sure

OpenStudy (mathmate):

No problem! :)